# GSQA: An End-to-End Model for Generative Spoken Question Answering

### 15 Dec 2023 v1

@@ 5.3 Performance at different WERs
Fig. 3: We sample the 8000 data within NMSQA-dev to verify the impact on cascade model under different Word Error Rates(WERs). Compared to cascade model, Unit-LongT5-TQA is more stable when higher WER leads to cascade model’s poor performance.
From the Fig.3, we can observe that Cascade model is quite sensitive to the impact of ASR WER. When ASR WER is low, we can see that cascade model performs well, which can get  50 F1-score. However, when the ASR WER increases, the drop of cascade model’s performance is obvious, dropping from 50 to 40 F1-score. On the other hand, from Table1, though our model (Unit-LongT5-Unit) performance is not as good as cascade model’s on average, we can see that our model can be more stable when ASR WER increases, even has better performance on high WER than cascade model.
@@ 5.4 Models parameters
For a fair comparison, we strive to ensure the number of parameters across all models remains relatively consistent. Upon examination, the cascade approach is found to have a larger number of parameters. This is primarily because it requires stacking various models together. Since the performance of each model should not be compromised, opting for smaller models is not feasible, leading to a larger overall parameter size. In contrast, end-to-end models rely on a single model to accomplish the task, resulting in a reduced number of parameters. We observed that for generative QA, there isn’t a stringent requirement for an exceptionally powerful speech-to-unit model. It’s sufficient to encode the content’s information into discrete units. Instead, a more potent QA model is needed to infer the answers to the questions. Detailed parameter counts are presented in Table 2. It is evident that our model has the fewest parameters among all methods. Notably, our model outperforms the others across various metrics.
@@ 6 Conclusion
In conclusion, we introduced the first end-to-end Generative Spoken Question Answering (GSQA) model, bridging the gap between extractive and abstractive QA tasks. By combining the advantages of cascade and textless models, our GSQA model demonstrates superior performance on extractive QA datasets, outperforming extractive spoken model. Furthermore, in the challenging abstractive zero-shot domain, our model exhibits competitive capabilities. We also highlighted the importance of pre-training the LM on textual data to enhance the unit-to-unit model’s semantic understanding. Future research may focus on improving the model’s performance with human narration and exploring strategies to enhance generalization across diverse spoken QA tasks.

### 25 Dec 2023 v2

@@ 5.2 Performance at different WERs
Fig. 3: We sample the 8000 data within NMSQA-dev to verify the impact on the cascaded model under different Word Error Rates (WERs) with different ASR systems.
Fig. 3 illustrates the superior stability and efficiency of our E2E model for SQA, especially compared to the cascaded model’s sensitivity to ASR WERs. The cascaded model, using ASR systems Whisper-medium.en and Whisper-small.en, demonstrates marked performance degradation with increasing WERs, whereas our E2E model maintains consistency even in high WER scenarios. Additionally, the cascaded model, particularly with Whisper-medium.en, requires significantly larger parameters, escalating computational resource demands. While enhancing the ASR system improves cascaded model performance, this approach requires extensive paired text-speech data, larger model scales, and considerable training costs for both LLMs and the TTS system, making it extremely expensive. In contrast, our smaller E2E model potentially offers comparable performance with greater training efficiency. Furthermore, the E2E model circumvents the limitations of languages without written forms, which are insurmountable for ASR-based systems. These advantages firmly position our E2E model as a more stable, resource-efficient, and universally applicable solution for SQA tasks.
@@ 5.3 Models parameters
For a fair comparison, we strive to ensure the number of parameters across all models remains relatively consistent. Upon examination, the cascaded method is found to have a large number of parameters. This is primarily because it requires stacking three huge models. Since the performance of each model should not be compromised, opting for smaller models is not feasible, leading to a larger overall parameter size. In contrast, end-to-end models rely on a single model to accomplish SQA tasks, resulting in a reduced number of parameters. We observed that for abstractive QA, there isn’t a stringent requirement for a compelling speech-to-unit model. It is sufficient to encode the content’s information into discrete units. Instead, a more potent QA model is needed to infer the answers to the questions. Detailed parameter counts are presented in Table 2. Our model has the fewest parameters among all models. Notably, our model outperforms the others across various metrics.
@@ 6 Conclusion
In conclusion, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model, bridging the gap between spoken extractive and abstractive QA tasks. By combining the advantages of cascaded and textless models, our GSQA model demonstrates superior performance on extractive QA datasets, outperforming the previous E2E extractive SQA models. Furthermore, in the challenging abstractive zero-shot domain, our model exhibits competitive capabilities. We also highlighted the importance of pretraining the language model on textual data to enhance the unit-to-unit model’s capability of semantic understanding. Our future research would concentrate on improving the model’s performance with human narrations and exploring training strategies to enhance generalization ability across diverse SQA tasks.